# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 6  Computing all reduced dominators
We now demonstrate an immediate application of Observation˜1 and our data structures, which we will frequently use to compute the reduced dominators:
Lemma 13.
Let $g\in P$ be a fixed point, and let $s,m,t\in\partial P$ with $m\in[s,t]$. Suppose $g$ lies in $\mathcal{E}[s,t]$. Then there exists a unique maximal chain $u,v]\subset[s,t]$ that is visible to $g$ and contains $m$, which can be found in $O(\log n)$ time.
Proof.
@@
Consider the polygon $P^{\prime}$ bounded by $S_{s}$, $[s,t]$, and $S_{t}$. Since $g$ sees $m$, the segment $\overline{g\,m}$ splits $P^{\prime}$ into two polygons, $P_{s}\subset P$ and $P_{t}\subset P$, containing $s$ and $t$ respectively (see Figure˜9). Let $u^{*}$ be the maximal point in $[s,m]$ such that $\overline{g\,u^{*}}$ intersects $S_{s}$ in more than one point (i.e., not only in $g$). We first show that $g$ sees all of $u^{*},m]$. Indeed, consider the polygon $P_{u^{*}}$ bounded by $u^{*},m]$, $\overline{m\,g}$, and $\overline{g\,u^{*}}$. Because $g\in\mathcal{E}[s,t]$, it lies left of all edges of $P_{u^{*}}$ and thus, by Observation˜1, $g$ sees all of $u^{*},m]$. By construction, $P_{u^{*}}$ contains no point of $S_{s}$ in its interior and hence lies within $P_{s}\subset P$. Therefore, $g$ indeed sees $u^{*},m]$ in $P$.
If $s\neq u^{*}$, then $g$ does not see any point $u^{\prime}\in[s,u^{*})$, since by definition of $u^{*}$, the segment $\overline{g\,u^{\prime}}$ intersects an edge of $S_{s}$ in its interior. It remains to show that $u^{*}=u$. Let $r^{*}$ be the ray from $g$ through $u^{*}$. We claim that $r^{*}=r_{s}$. Indeed, since $\overline{g\,u^{*}}$ intersects $S_{s}$, if this intersection occurs at a point $r$ not lying on the last edge of $S_{s}$, then $S_{s}$ could be shortened by the segment $\overline{r\,g}\subset\overline{u^{*}\,g}\subset P$. This contradicts the definition of $S_{s}$ as a shortest path. Hence $r^{*}=r_{s}$, and we conclude that $u^{*}=u$. By symmetry, the same argument applies to $r_{t}$, yielding the maximal visible chain $[u,v]$. ∎
@@ 6.1  A linear-size set $\mathcal{D}^{\prime}$ that contains all reduced good dominators in $\mathcal{D}$
We compute a linear-size set $\mathcal{D}^{\prime}$ that contains all guards in the set of reduced good dominators $\mathcal{D}$ (Definition˜6) via a sliding window. Formally, we maintain a pair of indices $(i,j)$ subject to incrementing $i$ and incrementing $j$. During this procedure, we maintain $\mathcal{E}[i-1,j+1]$ (and $\mathcal{E}[i-1,j+2]$) via Sliding-Window data structure (3) and we prove that this data structure can answer an advanced query:
Lemma 14.
Given vertices $i$ and $j$ with $i\leq j$, and $\mathcal{E}[i-1,j+1]$ in Sliding-Window data structure (3), we can compute the unique good dominator for $(i,j)$ (see Definition˜3), or return that no good dominator exists for $(i,j)$, in $O(\log n)$ time.
Proof.
@@
We apply Lemma˜13 to compute the maximal interval $u_{\text{max}},v_{\text{max}}]$ that is visible to $g^{*}$ (or $g^{\prime}$) that includes $i$ in logarithmic time. If $i,j]\subset[u_{\text{max}},v_{\text{max}}]$ then we return this guard as the segment-intersecting or angle-maximizing dominator for $i,j$. Otherwise, this procedure certified that no such dominator exists. ∎
Lemma 15.
@@
Suppose otherwise that $j^{\prime}$ lies left of $\overline{i\,j}$. Then because $P$ lies in general position, $j^{\prime}$ must lie strictly left of this line. However, then the interior of $\overline{i\,j^{\prime}}$ lies outside of the polygon $P^{\prime}$ defined by $\overline{i\,j}$ and $[i,j]$. We observed that $\mathcal{E}[i-1,j+1]$ is contained in $P^{\prime}$ and since $\mathcal{E}[i-1,j^{\prime}+1]\subset\mathcal{E}[i-1,j+1]$ it follows that no segment-intersecting dominator ($\mathbf{A}$) exists for $(i,j^{\prime})$. Finally if $j^{\prime}$ lies left of $\overline{i\,j}$ and $\mathcal{E}[i-1,j^{\prime}+1]$ is strictly left $\overline{i\,j^{\prime}}$ then no point in $\mathcal{E}[i-1,j^{\prime}+1]$ can see $j^{\prime}$ (see Figure˜12), Thus, there also exists no angle-maximizing dominator ($\mathbf{B}$) for $(i,j^{\prime})$.
Case 3: the potential dominator $g$ exists for $(i,j)$, but its maximal visible chain $u_{\text{max}},v_{\text{max}}]$ satisfies $u_{\text{max}}\notin[i-1,i]$ or $v_{\text{max}}\notin[j,j+1]$. We reduce this case to Case 2. Let $P^{\uparrow}$ denote the subpolygon of $\mathcal{E}[i-1,j+1]$ left of $\overline{i\,j}$. Consider any point $p\in P^{\uparrow}$. If $p$ sees both $i$ and $j$, then $(p,[i,j])$ forms a good guard (it satisfies Condition II). By Lemma˜5, there then exists a good dominator for $(i,j)$ that dominates $(p,[i,j])$, contradicting the assumption of the lemma. Hence, no point in $P^{\uparrow}$ sees both $i$ and $j$. Since any potential dominator for $(i,j^{\prime})$ with $j^{\prime}>j$ must see at least $[i,j]$, we may regard the visibility core $\mathcal{E}[i-1,j+1]$ as excluding $P^{\uparrow}$. This adjusted visibility core now lies strictly right of $\overline{i\,j}$ and thus has no potential dominator. Consequently, the situation reduces to Case 2. ∎
@@
Our strategy.
Let $(i,j)$ be the current index pair. We apply Lemma˜14 to find out whether a good dominator $(g,[u_{\text{max}},v_{\text{max}}])$ exists for $(i,j)$. If it exists, then Lemma˜14 returns it and we add it to $\mathcal{D}^{\prime}$. We then increment either $i$ or $j$ via the following strategy:
@@
Correctness.
We prove that we maintain our invariant. If we ever observe a pair $(i,j)$ for which there exists no good dominator, then by Lemma˜15 we note that there exists no good dominator for $(i,j^{\prime})$ with $j^{\prime}\geq i$ and we may safely increment $i$ in this case.
@@
If there exists a dominator $(g,[u_{\text{max}},v_{\text{max}}])$ for $(i,j)$ where $v_{\text{max}}<j+1$, then we check whether there exists a good dominator for $(i,j+1)$. If so, then any good dominator for $(i^{\prime},j)$ with $i<i^{\prime}<j$ will be dominated by this dominator and the invariant is thus maintained when increasing $j$. If not, then the previous argument implies that there is also no good dominator for $j^{\prime}>j+1$. It follows that the invariant is maintained after incrementing $i$. ∎
@@ 6.2  Computing a conforming sliding sequence $\sigma$
If $s=t$ then we simply output the guard $(s,[s,t])$. Otherwise, define $i^{*}$ as $i$ if $s<i$ and as $i+1$ if $s=i$. Define $j^{*}$ similarly as $j$ if $t>j$ and as j1j-1 if $t=j$. If $i^{*}=j^{*}$ then $i^{*}$ can see both $s$ and $t$ and we output the guard $(i^{*},[s,t])$. Otherwise, we know that $i^{*}<j^{*}$. We compute a constant-size set of candidate points $\mathbb{C}\subset\mathcal{E}[i-1,j+1]$.
@@
Computing the candidate set $\mathbb{C}$. We deploy a very similar strategy to Lemma˜5. We perform an emptiness query on $\mathcal{E}[i-1,j+1]$ which, if the visibility core is non-empty, returns a point $\gamma$. If the visibility core is empty we output that there exists no guard that guards $[s,t]$. We perform two ray shooting queries using the supporting ray from $i^{*}$ to $j^{*}$, and the ray from $j^{*}$ to $i^{*}$, in $O(\log n)$ time. If both rays do not hit $\mathcal{E}[i-1,j+1]$ then $\mathcal{E}[i-1,j+1]$ lies either strictly right or strictly left of $\overline{i^{*}\,j^{*}}$. We compare $\gamma$ to this supporting line to test in constant time whether the visibility core $\mathcal{E}[i-1,j+1]$ is strictly left of $\overline{i^{*}\,j^{*}}$. If so, then we define $g^{*}$ as the point in $\mathcal{E}[i-1,j+1]$ that maximizes the angle $\sphericalangle(i^{*},g^{*},j^{*})$, which we can compute in $O(\log n)$ time using an angle maximizing query. We can also use both ray shooting queries to detect whether the interior of $\overline{i^{*}\,j^{*}}$ intersects $\mathcal{E}[i-1,j+1]$, and if it does, we define $g^{\prime}$ as the last point of intersection of the open directed segment $\overline{i^{*}\,j^{*}}$ and $\mathcal{E}[i-1,j+1]$.
@@
Finally, we compute the shortest path $S(s,t)$ from $s$ to $t$ and test whether $S(s,t)$ is left-turning. If $S(s,t)$ is left-turning then we take the supporting halflines $\ell_{s}$ and $\ell_{t}$ of the first and last edge of $S(s,t)$ (these lines are directed away from $s$ and $t$, respectively). We compute $g_{1}$ and $g_{2}$: the respective last points of intersection of these halflines with $\mathcal{E}[i-1,j+1]$ in $O(\log n)$ time. The set $\mathbb{C}$ is formed by $\gamma,g^{*},g^{\prime},g_{1}$ and $g_{2}$ (if the respective point exists). For each $g\in\mathbb{C}$, we observe that $g\in\mathcal{E}[i-1,j+1]\subset\mathcal{E}[s,t]$ and so we invoke Lemma˜13 to compute whether $g$ can see $[s,t]$ in $O(\log n)$ time. If none of the points in $\mathbb{C}$ has this property, we output that no guard $(g,[s,t])$ exists.
@@
Correctness. Suppose that no good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists. Then if any guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists it must be that $\sphericalangle(s,g,t)>\pi$. This in turn implies that $S(s,t)$ is a left-turning chain, and so a point $p^{\prime}\in\mathcal{E}[i-1,j+1]\subset\mathcal{E}[s,t]$ can see $[u,v]$ if and only if $p^{\prime}$ lies right of $\ell_{s}$ and left of $\ell_{t}$. If there exists at least one point $p^{\prime}\in\mathcal{E}[i-1,j+1]$ right of $\ell_{s}$ and left of $\ell_{t}$, then one of $\{\gamma,g_{1},g_{2}\}$ exists and lies right of $\ell_{s}$ and left of $\ell_{t}$. It follows that a guard $(g,[s,t])$ exists if and only if our algorithm outputs such a guard. Suppose otherwise that a good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists. Then $\sphericalangle(i^{*},g,j^{*})\leq\pi$. We make a case distinction.
@@
Case 1: $g^{*}$ does not exist. Via an identical argument as in Lemma˜5, we then get that $\mathcal{E}[i-1,j+1]$ must intersect the interior of $\overline{i^{*}\,j^{*}}$. In this case, any point of intersection between $\mathcal{E}[i-1,j-1]$ and $\overline{i^{*}\,j^{*}}$, and in particular our precomputed point $g^{\prime}$, can see $[s,t]$.
@@
Case 2: $g^{*}$ exists. If $g^{*}$ exists, then the visibility core $\mathcal{E}[i-1,j+1]$ is strictly left of $\overline{i^{*}\,j^{*}}$ by definition, and thus $\sphericalangle(i^{*},g^{*},j^{*})<\pi$. We now define the polygon $P_{g}$ that is formed by $[s,t]$ and $\overline{g\,s}$ and $\overline{t\,g}$. The point $g^{*}$ by definition lies in $\mathcal{E}[i-1,j+1]\subset\mathcal{E}[s,t]$. If $g^{*}$ is in $P_{g}$ then it lies in $P_{g}\cap\mathcal{E}[s,t]$ and thus sees all of $[s,t]$. We now have an identical construction to Case 2 of Lemma˜5 and conclude that $g^{*}$ can see $[s,t]$.
@@
Conclusion. If a good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists then either $g^{\prime}$ or $g^{*}$ can see $[s,t]$. If no good guard with $g\in\mathcal{E}[i-1,j+1]$ exists, then any point $p\in\mathcal{E}[i-1,j+1]\subset\mathcal{E}[s,t]$ can see $[s,t]$ if and only if $p$ lies left of $\ell_{s}$ and right of $\ell_{t}$. So, if none of $\{\gamma,g^{*},g^{\prime},g_{1},g_{2}\}$ guard the chain $[s,t]$, then no guard in $\mathcal{E}[i-1,j+1]$ can. ∎
Theorem 18.
@@
Proof.
Recall that a sliding sequence has for consecutive pairs $((i,j),(i^{\prime},j^{\prime}))$ that $(i^{\prime}-i,j^{\prime}-j)\in\{(1,0),(0,1)\}$. Formally, we construct $\sigma$ by focusing on a different property. We define the discrete maximum function $d_{\max}$ that takes any integer $i\leq n$ and returns the maximum integer $j^{*}$ such that there exists a guard $(g,[i,j^{*}])$ with $g\in\mathcal{E}[i-1,j^{*}+1]$. Note that for any $j\in[i,d_{\max}(i)]$, the visibility core $\mathcal{E}[i-1,j^{*}+1]\subset\mathcal{E}[i-1,j+1]$ and so for all $j\in[i,d_{\max}(i)]$ there exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. We prove that any $(i,j)$ for which there exists a guard $(g,[u,\textnormal{{next}}(u)])$ with $u\in[i-1,i)$ and $\textnormal{{next}}(u)\in(j,j+1]$ is included in any sliding sequence $\sigma^{*}$ that includes all pairs $(i,j)$ with $j\in[i,d_{\max}(i)]$ where ‘no strictly better’ pair $(i^{\prime},j^{\prime})$ exists. Formally, we construct a sliding sequence $\sigma^{*}$ such that:
@@
We claim that $\sigma^{*}$ is conforming. Indeed, for fixed indices $(i,j)$, if there exists a guard $(g,[u,\textnormal{{next}}(u)])$ with $u\in[i-1,i)$ and $\textnormal{{next}}(u)\in(j,j+1]$, then there also exists a guard $(g^{\prime},[i,j])$ with $g^{\prime}\in\mathcal{E}[i-1,j+1]$ and so $j\in[i,d_{\max}(i)]$. Thus $(i,j)\in\sigma^{*}$ if there exists no strictly better pair $(i^{\prime},j^{\prime})$. Suppose for the sake of contradiction that there exist indices $i^{\prime}<i$ and $j^{\prime}>j$ with $j^{\prime}\in d_{\max}(i^{\prime})$. Then by definition, there exists a guard $(g^{\prime\prime},[i^{\prime},j^{\prime}])$ with $g^{\prime\prime}\in\mathcal{E}[i^{\prime}-1,j^{\prime}+1]\subset\mathcal{E}[i-1,j+1]$ but this contradicts the maximality of $\textnormal{{next}}(u)$ for our original guard $g$.
Computing $\sigma^{*}$.
We maintain $(i,j)$ subject to incrementing either $i$ or $j$ and maintain the following invariants:

### v4

@@ 6 Computing all reduced dominators
We now demonstrate an immediate application of Section˜2 and our data structures, which we will frequently use to compute the reduced dominators:
Lemma 6.1.
Let $g\in P$ be a fixed point, and let $s,m,t\in\partial P$ with $m\in[s,t]$. Suppose $g$ lies in $\mathcal{E}[s,t]$. Then there exists a unique maximal chain $u,v]\subseteq[s,t]$ that is visible to $g$ and contains $m$, which can be found in $O(\log n)$ time.
Proof 6.2.
@@
Consider the polygon $P^{\prime}$ bounded by $S_{s}$, $[s,t]$, and $S_{t}$. Since $g$ sees $m$, the segment $\overline{g\,m}$ splits $P^{\prime}$ into two polygons, $P_{s}\subseteq P$ and $P_{t}\subseteq P$, containing $s$ and $t$ respectively (see Figure˜9). Let $u^{*}$ be the maximal point in $[s,m]$ such that $\overline{g\,u^{*}}$ intersects $S_{s}$ in more than one point (i.e., not only in $g$). We first show that $g$ sees all of $u^{*},m]$. Indeed, consider the polygon $P_{u^{*}}$ bounded by $u^{*},m]$, $\overline{m\,g}$, and $\overline{g\,u^{*}}$. Because $g\in\mathcal{E}[s,t]$, it lies left of all edges of $P_{u^{*}}$ and thus, by Section˜2, $g$ sees all of $u^{*},m]$. By construction, $P_{u^{*}}$ contains no point of $S_{s}$ in its interior and hence lies within $P_{s}\subseteq P$. Therefore, $g$ indeed sees $u^{*},m]$ in $P$.
If $s\neq u^{*}$, then $g$ does not see any point $u^{\prime}\in[s,u^{*})$, since by definition of $u^{*}$, the segment $\overline{g\,u^{\prime}}$ intersects an edge of $S_{s}$ in its interior. It remains to show that $u^{*}=u$. Let $r^{*}$ be the ray from $g$ through $u^{*}$. We claim that $r^{*}=r_{s}$. Indeed, since $\overline{g\,u^{*}}$ intersects $S_{s}$, if this intersection occurs at a point $r$ not lying on the last edge of $S_{s}$, then $S_{s}$ could be shortened by the segment $\overline{r\,g}\subseteq\overline{u^{*}\,g}\subseteq P$. This contradicts the definition of $S_{s}$ as a shortest path. Hence $r^{*}=r_{s}$, and we conclude that $u^{*}=u$. By symmetry, the same argument applies to $r_{t}$, yielding the maximal visible chain $[u,v]$.
@@ 6.1 A linear-size set $\mathcal{D}^{\prime}$ that contains all reduced good dominators in $\mathcal{D}$
We compute a linear-size set $\mathcal{D}^{\prime}$ that contains all guards in the set of reduced good dominators $\mathcal{D}$ (Definition˜4.12) via a sliding window. Formally, we maintain a pair of indices $(i,j)$ subject to incrementing $i$ and incrementing $j$. During this procedure, we maintain $\mathcal{E}[i-1,j+1]$ (and $\mathcal{E}[i-1,j+2]$) via Sliding-Window data structure (3) and we prove that this data structure can answer an advanced query:
Lemma 6.3.
Given vertices $i$ and $j$ with $i\leq j$, and $\mathcal{E}[i-1,j+1]$ in Sliding-Window data structure (3), we can compute the unique good dominator for $(i,j)$ (see Definition˜4.7), or return that no good dominator exists for $(i,j)$, in $O(\log n)$ time.
Proof 6.4.
@@
We apply Lemma˜6.1 to compute the maximal interval $u_{\text{max}},v_{\text{max}}]$ that is visible to $g^{*}$ (or $g^{\prime}$) that includes $i$ in logarithmic time. If $i,j]\subseteq[u_{\text{max}},v_{\text{max}}]$ then we return this guard as the segment-intersecting or angle-maximizing dominator for $i,j$. Otherwise, this procedure certified that no such dominator exists.
Lemma 6.5.
@@
Suppose otherwise that $j^{\prime}$ lies left of $\overline{i\,j}$. Then because $P$ lies in general position, $j^{\prime}$ must lie strictly left of this line. However, then the interior of $\overline{i\,j^{\prime}}$ lies outside of the polygon $P^{\prime}$ defined by $\overline{i\,j}$ and $[i,j]$. We observed that $\mathcal{E}[i-1,j+1]$ is contained in $P^{\prime}$ and since $\mathcal{E}[i-1,j^{\prime}+1]\subseteq\mathcal{E}[i-1,j+1]$ it follows that no segment-intersecting dominator ($\mathbf{A}$) exists for $(i,j^{\prime})$. Finally if $j^{\prime}$ lies left of $\overline{i\,j}$ and $\mathcal{E}[i-1,j^{\prime}+1]$ is strictly left $\overline{i\,j^{\prime}}$ then no point in $\mathcal{E}[i-1,j^{\prime}+1]$ can see $j^{\prime}$ (see Figure˜12), Thus, there also exists no angle-maximizing dominator ($\mathbf{B}$) for $(i,j^{\prime})$.
Case 3: the potential dominator $g$ exists for $(i,j)$, but its maximal visible chain $u_{\text{max}},v_{\text{max}}]$ satisfies $u_{\text{max}}\notin[i-1,i]$ or $v_{\text{max}}\notin[j,j+1]$. We reduce this case to Case 2. Let $P^{\uparrow}$ denote the subpolygon of $\mathcal{E}[i-1,j+1]$ left of $\overline{i\,j}$. Consider any point $p\in P^{\uparrow}$. If $p$ sees both $i$ and $j$, then $(p,[i,j])$ forms a good guard (it satisfies Condition II). By Lemma˜4.9, there then exists a good dominator for $(i,j)$ that dominates $(p,[i,j])$, contradicting the assumption of the lemma. Hence, no point in $P^{\uparrow}$ sees both $i$ and $j$. Since any potential dominator for $(i,j^{\prime})$ with $j^{\prime}>j$ must see at least $[i,j]$, we may regard the visibility core $\mathcal{E}[i-1,j+1]$ as excluding $P^{\uparrow}$. This adjusted visibility core now lies strictly right of $\overline{i\,j}$ and thus has no potential dominator. Consequently, the situation reduces to Case 2.
@@
Our strategy.  Let $(i,j)$ be the current index pair. We apply Lemma˜6.3 to find out whether a good dominator $(g,[u_{\text{max}},v_{\text{max}}])$ exists for $(i,j)$. If it exists, then Lemma˜6.3 returns it and we add it to $\mathcal{D}^{\prime}$. We then increment either $i$ or $j$ via the following strategy:
@@
Correctness.  We prove that we maintain our invariant. If we ever observe a pair $(i,j)$ for which there exists no good dominator, then by Lemma˜6.5 we note that there exists no good dominator for $(i,j^{\prime})$ with $j^{\prime}\geq i$ and we may safely increment $i$ in this case.
@@
If there exists a dominator $(g,[u_{\text{max}},v_{\text{max}}])$ for $(i,j)$ where $v_{\text{max}}<j+1$, then we check whether there exists a good dominator for $(i,j+1)$. If so, then any good dominator for $(i^{\prime},j)$ with $i<i^{\prime}<j$ will be dominated by this dominator and the invariant is thus maintained when increasing $j$. If not, then the previous argument implies that there is also no good dominator for $j^{\prime}>j+1$. It follows that the invariant is maintained after incrementing $i$.
@@ 6.2 Computing a conforming sliding sequence $\sigma$
If $s=t$ then we simply output the guard $(s,[s,t])$. Otherwise, define $i^{*}$ as $i$ if $s<i$ and as $i+1$ if $s=i$. Define $j^{*}$ similarly as $j$ if $t>j$ and as j1j-1 if $t=j$. If $i^{*}=j^{*}$ then $i^{*}$ can see both $s$ and $t$ and we output the guard $(i^{*},[s,t])$. Otherwise, we know that $i^{*}<j^{*}$. We compute a constant-size set of candidate points $\mathbb{C}\subseteq\mathcal{E}[i-1,j+1]$.
@@
Computing the candidate set $\mathbb{C}$. We deploy a very similar strategy to Lemma˜4.9. We perform an emptiness query on $\mathcal{E}[i-1,j+1]$ which, if the visibility core is non-empty, returns a point $\gamma$. If the visibility core is empty we output that there exists no guard that guards $[s,t]$. We perform two ray shooting queries using the supporting ray from $i^{*}$ to $j^{*}$, and the ray from $j^{*}$ to $i^{*}$, in $O(\log n)$ time. If both rays do not hit $\mathcal{E}[i-1,j+1]$ then $\mathcal{E}[i-1,j+1]$ lies either strictly right or strictly left of $\overline{i^{*}\,j^{*}}$. We compare $\gamma$ to this supporting line to test in constant time whether the visibility core $\mathcal{E}[i-1,j+1]$ is strictly left of $\overline{i^{*}\,j^{*}}$. If so, then we define $g^{*}$ as the point in $\mathcal{E}[i-1,j+1]$ that maximizes the angle $\sphericalangle(i^{*},g^{*},j^{*})$, which we can compute in $O(\log n)$ time using an angle maximizing query. We can also use both ray shooting queries to detect whether the interior of $\overline{i^{*}\,j^{*}}$ intersects $\mathcal{E}[i-1,j+1]$, and if it does, we define $g^{\prime}$ as the last point of intersection of the open directed segment $\overline{i^{*}\,j^{*}}$ and $\mathcal{E}[i-1,j+1]$.
@@
Finally, we compute the shortest path $S(s,t)$ from $s$ to $t$ and test whether $S(s,t)$ is left-turning. If $S(s,t)$ is left-turning then we take the supporting halflines $\ell_{s}$ and $\ell_{t}$ of the first and last edge of $S(s,t)$ (these lines are directed away from $s$ and $t$, respectively). We compute $g_{1}$ and $g_{2}$: the respective last points of intersection of these halflines with $\mathcal{E}[i-1,j+1]$ in $O(\log n)$ time. The set $\mathbb{C}$ is formed by $\gamma,g^{*},g^{\prime},g_{1}$ and $g_{2}$ (if the respective point exists). For each $g\in\mathbb{C}$, we observe that $g\in\mathcal{E}[i-1,j+1]\subseteq\mathcal{E}[s,t]$ and so we invoke Lemma˜6.1 to compute whether $g$ can see $[s,t]$ in $O(\log n)$ time. If none of the points in $\mathbb{C}$ has this property, we output that no guard $(g,[s,t])$ exists.
@@
Correctness. Suppose that no good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists. Then if any guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists it must be that $\sphericalangle(s,g,t)>\pi$. This in turn implies that $S(s,t)$ is a left-turning chain, and so a point $p^{\prime}\in\mathcal{E}[i-1,j+1]\subseteq\mathcal{E}[s,t]$ can see $[u,v]$ if and only if $p^{\prime}$ lies right of $\ell_{s}$ and left of $\ell_{t}$. If there exists at least one point $p^{\prime}\in\mathcal{E}[i-1,j+1]$ right of $\ell_{s}$ and left of $\ell_{t}$, then one of $\{\gamma,g_{1},g_{2}\}$ exists and lies right of $\ell_{s}$ and left of $\ell_{t}$. It follows that a guard $(g,[s,t])$ exists if and only if our algorithm outputs such a guard. Suppose otherwise that a good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists. Then $\sphericalangle(i^{*},g,j^{*})\leq\pi$. We make a case distinction.
@@
Case 1: $g^{*}$ does not exist. Via an identical argument as in Lemma˜4.9, we then get that $\mathcal{E}[i-1,j+1]$ must intersect the interior of $\overline{i^{*}\,j^{*}}$. In this case, any point of intersection between $\mathcal{E}[i-1,j-1]$ and $\overline{i^{*}\,j^{*}}$, and in particular our precomputed point $g^{\prime}$, can see $[s,t]$.
@@
Case 2: $g^{*}$ exists. If $g^{*}$ exists, then the visibility core $\mathcal{E}[i-1,j+1]$ is strictly left of $\overline{i^{*}\,j^{*}}$ by definition, and thus $\sphericalangle(i^{*},g^{*},j^{*})<\pi$. We now define the polygon $P_{g}$ that is formed by $[s,t]$ and $\overline{g\,s}$ and $\overline{t\,g}$. The point $g^{*}$ by definition lies in $\mathcal{E}[i-1,j+1]\subseteq\mathcal{E}[s,t]$. If $g^{*}$ is in $P_{g}$ then it lies in $P_{g}\cap\mathcal{E}[s,t]$ and thus sees all of $[s,t]$. We now have an identical construction to Case 2 of Lemma˜4.9 and conclude that $g^{*}$ can see $[s,t]$.
@@
Conclusion. If a good guard $(g,[s,t])$ with $g\in\mathcal{E}[i-1,j+1]$ exists then either $g^{\prime}$ or $g^{*}$ can see $[s,t]$. If no good guard with $g\in\mathcal{E}[i-1,j+1]$ exists, then any point $p\in\mathcal{E}[i-1,j+1]\subseteq\mathcal{E}[s,t]$ can see $[s,t]$ if and only if $p$ lies left of $\ell_{s}$ and right of $\ell_{t}$. So, if none of $\{\gamma,g^{*},g^{\prime},g_{1},g_{2}\}$ guard the chain $[s,t]$, then no guard in $\mathcal{E}[i-1,j+1]$ can.
Theorem 6.11.
@@
Proof 6.12.
Recall that a sliding sequence has for consecutive pairs $((i,j),(i^{\prime},j^{\prime}))$ that $(i^{\prime}-i,j^{\prime}-j)\in\{(1,0),(0,1)\}$. Formally, we construct $\sigma$ by focusing on a different property. We define the discrete maximum function $d_{\max}$ that takes any integer $i\leq n$ and returns the maximum integer $j^{*}$ such that there exists a guard $(g,[i,j^{*}])$ with $g\in\mathcal{E}[i-1,j^{*}+1]$. Note that for any $j\in[i,d_{\max}(i)]$, the visibility core $\mathcal{E}[i-1,j^{*}+1]\subseteq\mathcal{E}[i-1,j+1]$ and so for all $j\in[i,d_{\max}(i)]$ there exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. We prove that any $(i,j)$ for which there exists a guard $(g,[u,\textnormal{{next}}(u)])$ with $u\in[i-1,i)$ and $\textnormal{{next}}(u)\in(j,j+1]$ is included in any sliding sequence $\sigma^{*}$ that includes all pairs $(i,j)$ with $j\in[i,d_{\max}(i)]$ where ‘no strictly better’ pair $(i^{\prime},j^{\prime})$ exists. Formally, we construct a sliding sequence $\sigma^{*}$ such that:
@@
We claim that $\sigma^{*}$ is conforming. Indeed, for fixed indices $(i,j)$, if there exists a guard $(g,[u,\textnormal{{next}}(u)])$ with $u\in[i-1,i)$ and $\textnormal{{next}}(u)\in(j,j+1]$, then there also exists a guard $(g^{\prime},[i,j])$ with $g^{\prime}\in\mathcal{E}[i-1,j+1]$ and so $j\in[i,d_{\max}(i)]$. Thus $(i,j)\in\sigma^{*}$ if there exists no strictly better pair $(i^{\prime},j^{\prime})$. Suppose for the sake of contradiction that there exist indices $i^{\prime}<i$ and $j^{\prime}>j$ with $j^{\prime}\in d_{\max}(i^{\prime})$. Then by definition, there exists a guard $(g^{\prime\prime},[i^{\prime},j^{\prime}])$ with $g^{\prime\prime}\in\mathcal{E}[i^{\prime}-1,j^{\prime}+1]\subseteq\mathcal{E}[i-1,j+1]$ but this contradicts the maximality of $\textnormal{{next}}(u)$ for our original guard $g$.
Computing $\sigma^{*}$.  We maintain $(i,j)$ subject to incrementing either $i$ or $j$ and maintain the following invariants: